An AC source has an output rms voltage of 71.0 V at a frequency of 41.5 Hz. The source is connected across a 25.5-mH inductor. (a) Find the inductive reactance of the circuit. (b) Find the rms current in the circuit. (c) Find the maximum current in the circuit.
Answer #1

Inductive reactance at those conditions is
XL = 2πfL = 2*3.14*41.5*25.5*10-3 = 6.645 ohms

I = E/X = 71/6.645 = 10.684 amps

the maximum current depends upon the circuit Whether they are in series or parallel.